CO2 will be produced when 1.60 mol of CS2 burns with 5.60 mol of O2. Determine the limiting reactant


1
Expert's answer
2021-02-28T06:19:38-0500

Solution:

The balanced chemical equation:

CS2 + 3O2 → CO2 + 2SO2

According to the equation above: n(CS2) = n(O2)/3


Moles of CS2 = 1.60 mol

Moles of O2 = 5.60 mol


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose CS2:

n(O2) = 3 × n(CS2) = 3 × 1.60 mol = 4.80 mol


The calculation above means that we need 4.80 mol of O2 to completely react with CS2.

We have 5.60 mol of O2 and therefore more than enough oxygen.

Thus oxygen (O2) is in excess and carbon disulfide (CS2) must be the limiting reagent.

Hence, the limiting reactant is carbon disulfide (CS2).


Answer: The limiting reactant is CS2.